Solve for the value of x

2x(0.4-0.3)=x(0.3-0.5)-0.2(x-1)

Answer by adamchapman(301)

2x(0.4-0.3)=x(0.3-0.5)-0.2(x-1)
2x(0.1)=x(-0.2)-0.2(x-1)
0.2x=-0.2x-0.2x-0.2
0.6x=-0.2
x=-1/3
I hope this helps,
Adam
P.S. please visit my website, it may be helpful to you. The address is www.geocities.com/quibowibbler